OpenStudy (anonymous):

How do you balance Cu3(PO4)2 + AgNO3 =

OpenStudy (anonymous):

first, determine what the products of the reaction will be

OpenStudy (anonymous):

so in this case, most likely you will have a double replacement where the Cu cation switches anions with the Ag cation. Determine each ion's charges

OpenStudy (anonymous):

silver is +1, NO3 is -1, PO4 is always a -3, and Cu must be +2 because it takes 3 Cu ions and 2 PO4 ions to make the entire molecule neutral. (2*-3+3*2=-6+6=0

OpenStudy (anonymous):

then, rearrange the molecules, replacing the Cu3(PO4)2 anion with the AgNO3 anion. this produces Cu(NO3)2 and Ag3PO4

OpenStudy (anonymous):

from there, list out your molecules in the equation and change coefficients until there are equal numbers of the same atom on either side of the equation.

OpenStudy (anonymous):

adjust until you end with matching numbers of atoms/ions

OpenStudy (anonymous):

looks like this: Cu3(PO4)2+AGNO3 ==> Cu(NO3)2+Ag3PO4 Reactants Products Cu-3 Cu-1 PO4-2 PO4-1 Ag-1 Ag-3 NO3-1 NO3-1

OpenStudy (anonymous):

I still do not get it, you have on one side 3 Cu and on the other side only 1 Cu, I thought you have to balance it so the one side has the same number as the other ???

OpenStudy (chmvijay):

2Cu3(PO4)2 + 6AgNO3 = 2Ag3(PO4) + 6CuNO3

OpenStudy (chmvijay):

did u get now @kajal

OpenStudy (anonymous):

no because you have on one side 2X(PO4)2 =4, and on the other side only 2(PO4) ,

OpenStudy (anonymous):

2Cu3(PO4)2 + 6AgNO3 = 6CuNO3+2Ag3(PO4)2 maybe???

OpenStudy (anonymous):

Cu3(PO4)2 + 3 AgNO3 = 3 CuNO3 + Ag3(PO4)2 or this?

OpenStudy (chmvijay):

ur right lol :(